OpenStudy (loser66):

Someone helps me please. What is the relationship between Ax =0 and Ax=b

OpenStudy (adi3):

I think b = 0

OpenStudy (adi3):

okay

OpenStudy (adi3):

why did u ask

OpenStudy (loser66):

Let r>0 \(f(x)=\begin{cases} x^rsin(1/x) ~~~~if x>0\\0~~~if x\leq 0\end{cases}\) how to prove f(x) differentiable?

OpenStudy (loser66):

@freckles

OpenStudy (freckles):

Well we need first establish that f is continuous at x=0 and then we need to check out that whole smoothness thing.

OpenStudy (freckles):

\[f(0)=0 \text{ is easy \to see now we just need \to \check if that limit thing is 0 } \\ \lim_{x \rightarrow 0^+} x^r \sin(\frac{1}{x}) =\lim_{x \rightarrow 0^-}0\] so you know that left limit at x=0 is also obviously 0 we just to make sure the other side is also 0 then we can move on to that smoothness requirement

OpenStudy (freckles):

\[-1 \le \sin(\frac{1}{x}) \le 1 \] I suggest using squeeze theorem

OpenStudy (fibonaccichick666):

Still need me, or you got it?

OpenStudy (loser66):

how? when left side is -1 and right side is 1

OpenStudy (freckles):

multiply both sides by x^r

OpenStudy (freckles):

all sides*

OpenStudy (loser66):

I need a lot of help. Appreciate any help

OpenStudy (freckles):

\[-x^r \le x^r \sin(\frac{1}{x}) \le x^r \]

OpenStudy (freckles):

now we know r>0 so you know both -x^r and x^r go to what as x goes to zero?

OpenStudy (fibonaccichick666):

not quite how squeeze is applied there, but yes, you need to use the periodic nature of sine. Then squeeze if you have already proved it in class

OpenStudy (loser66):

0

OpenStudy (freckles):

by the way the direction of the inequality isn't effected by our multiplication because x>0 since we are only looking at x goes to infinity so you just did the first part you showed that the function is continuous at x=0

OpenStudy (loser66):

so for any value of r, we have f'(0) exist, right?

OpenStudy (freckles):

now we need to check the smoothness at x=0 (I thought we were looking at r>0)

OpenStudy (loser66):

ok, that part is ok. Now another question: :) for which value of r, f'(x) continuous at 0?

OpenStudy (freckles):

well you know we haven't finished your first problem right?

OpenStudy (freckles):

anyways...for the second question you have here you need to find f'(x)

OpenStudy (freckles):

do that by finding it in pieces first

OpenStudy (freckles):

like find the derivative of both left and right sides of the function at x=0

OpenStudy (loser66):